OpenStudy (learningallday):

Which equation demonstrates the opposite of a sum property? A. (22 + 10) + 3 = (10 + 22) + 3 B. –5 + 13 = 13 + (–5) C. (12 + (–4)) + 14 = 12 + ((–4) + 14) D. –(–6 + 17) = 6 + (–17)

OpenStudy (anonymous):

the answer is C.

OpenStudy (learningallday):

thanks

OpenStudy (anonymous):

your welcome

OpenStudy (learningallday):

i got another one

OpenStudy (anonymous):

alrighty

OpenStudy (learningallday):

What value of z makes the statement true? –(z + (–17)) = 12 + 17 A. –46 B. –12 C. 12 D. 46

OpenStudy (anonymous):

i do believe A

OpenStudy (learningallday):

ok

OpenStudy (anonymous):

;-)

OpenStudy (learningallday):

–(5 + (–11.3)) + (–4.3) A. –20.6 B. –10.6 C. –2 D. 2

OpenStudy (learningallday):

:)

OpenStudy (anonymous):

D.

OpenStudy (learningallday):

What value of k makes the statement true? k + (–4.1) = –9.7

OpenStudy (learningallday):

:-) :) :) :) :) :) :) :) :) :) :)

OpenStudy (anonymous):

-5.6

OpenStudy (anonymous):

so do you have a kik

OpenStudy (learningallday):

yup

OpenStudy (anonymous):

may i add you

OpenStudy (anonymous):

im 16 btw

OpenStudy (anonymous):

i couldnt find it

OpenStudy (anonymous):

done

OpenStudy (learningallday):

What value of c makes the statement true? c – (–17.6) = –4 A. –21.6 B. –12.4 C. 12.4 D. 21.6

OpenStudy (anonymous):

A.